OpenStudy (anonymous):

what is 3x-2y=11 in slope intercept form

OpenStudy (anonymous):

3x-2y=11 Try to take this to form y=mx+c

OpenStudy (anonymous):

evrytime i do i get decimals

OpenStudy (anonymous):

Keep them in fractions

OpenStudy (anonymous):

my answer was y=-1/2+5.5

OpenStudy (anonymous):

y=-3/2-11/2

OpenStudy (anonymous):

3x-2y=11 2y=3x-11 y=3/2 x-11/2

OpenStudy (anonymous):

U missed the x

OpenStudy (anonymous):

sorry y=3/2x-11/2

OpenStudy (anonymous):

so does that mean for solveing the solution that you cant do it?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

how would you graph 11/2 on the y axis

OpenStudy (anonymous):

11/2=5.5 between 5 and 6

OpenStudy (anonymous):

alright thank you

OpenStudy (anonymous):

welcome